OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively.

OpenStudy (anonymous):

we need to know the first term and the common ratio

OpenStudy (anonymous):

let a be the first term and r be the common ratio then we have ar^4= -2 and ar^7=16 thus dividing the two eq. we hve r^3= 16/(-2)= -8 hence r =-2 now a(-2)^4= -2 or 16 a= -2 or a= -2/16 a=-1/8 so a_n= (-1/8)*(-2)^(n-1)

OpenStudy (imstuck):

If a is the first term, and r is the ratio, it looks to me like you are using the fact that the second term a_2 = -2. Wouldn't your identity, then, be -2 = a(r)^2-1 which is -2 = a(r)^1? I'm just confused as to where you get the exponent on the r in each one.

OpenStudy (mosaic):

The nth term of a geometric sequence is \(a_n = ar^{n-1}\) \( a_2 = ar \\ a_5 = ar^4 \\ a_5/a_2 = r^3 = 16/(-2) = -8 \\ r = -2 \\ a_2 = ar = -2 \\ a = -2/r = -2/-2 = 1 \\ a_n = (-2)^{n-1} \)

OpenStudy (imstuck):

Gotcha, mosaic. Thank you for that explanation...whoa

OpenStudy (mosaic):

You are welcome.

OpenStudy (anonymous):

Thank you!

OpenStudy (mosaic):

np.